# Zero-shot Safety Prediction for Autonomous Robots with Foundation World Models

### v1

@@ I Introduction
In the past, researchers primarily focused on training better controllers with world models, with insufficient attention given to how accurate the world of world models is — a crucial aspect for safety-critical robots. Typically, a world model does not directly predict observations, such as images from cameras or other sensor data, due to the constraints on time and computational resources. Instead, as shown on the left of Fig. 2, it extracts useful features into latent representations with an encoder, which is part of a Variational Autoencoder (VAE) [6], and then predicts based on these representations. Usually, the performance of the world model is evaluated with the Mean Square Error (MSE) between the predicted observation and ground truth; this metric considers only the aggregate performance of the reconstruction — and lacks the examination of the fine-grained and meaningful details of the prediction. This coarse checking does not distinguish slight but critical differences (e.g., if a car’s position shifts a few pixels, it may lead to a collision) that may cause devastating outcomes in safety predictions. Also, the latent representations do not have a physical meaning in standard models and cannot be used to evaluate whether a predicted latent state is safe or has other important characteristics. As a result, it becomes necessary to develop an additional classifier to check such critical aspects, akin to a safety check as discussed in our earlier work [5], which requires more data and adds noise.
@@
Our evaluation on two common simulated benchmarks demonstrates the value of foundation world models. We experiment with several baseline approaches, evaluation metrics, and different Large Language Models for latent prediction.
@@
    A segmentation-based metric for the accuracy of the surrogate dynamics by comparing how much each object in the observation deviated.
@@
    An experimental study of safety prediction where foundation world models show better performance despite not using any data compared to the existing world model and supervised learning methods.
Sec. II introduces the details of world models and our problem description. Sec. III and Sec. IV describe the approach and the results of the experiment. In Sec. V and Sec. VI, we will review the related work and discuss the conclusion and future work.
@@ II-A Observation Prediction with World Models
There is no single agreed-upon metric to evaluate the performance of a world model. In applications of world models, the controller-training tasks are usually judged on how well the resulting controller performs compared with traditional methods [1, 3] in terms of reward. Traditionally, evaluation metrics like MSE are commonly adopted as the standard measurement criteria for quantifying the predictive capacity of world models. Accordingly, we compute the MSE of the world model’s predicted observation in a pixel-wise manner:
@@ II-B Object-based Prediction Metric
Our insight, unused in earlier world models, is that the observation can be split into multiple meaningful objects by some segmentation algorithm $\Omega$: $\Omega(\textit{{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...]$. Each $\bm{\omega}$ is a segmentation matrix with $True$ and $False$ and has the same size of $\bm{y}$ as shown in Fig 3. The $True$ elements reflect the position of the segmented object. These objects in the observation are semantically important elements of the world that are assumed to have a causal relation to the future positions. The prediction of the objects’ positions will evaluated separately, unlike e.g. the image-wise MSE that takes all pixels into account and is affected by inconsequential entities like static objects/background that occupy most of the observation frame.
To tailor the evaluation to each object $\bm{\omega}$, we will calculate its centroid $\bm{\omega}^{c}=(\bm{\omega}^{c}_{x},\bm{\omega}^{c}_{y})$ (i.e., the center of mass) as:
@@
This process is described as a function CentrExt in Algorithm 1. Each prediction $\hat{\bm{\omega}}$ of a true object $\bm{\omega}$ will be quantified using the centroid distance (CD) defined with some norm $\mathcal{L}$, which can be, e.g., an L1 or L2-norm:
@@
The image errors between prediction and ground truth are compound, which contains a vector of CDs: $(\text{CD}(\bm{\omega}_{1},\hat{\bm{\omega}_{1}}),\text{CD}(\bm{\omega}_{2},\hat{\bm{\omega}_{2}}),...,\text{CD}(\bm{\omega}_{n},\hat{\bm{\omega}_{n}}))$ and can precisely reflect how well each part is predicted instead of providing all errors of unrelated pixels.
@@ II-C Safety Prediction Problem
$\varphi$ is a function of state $\bm{x}_{t}$ and shows the safety of the system at time t: $\varphi(\bm{x}_{t})$. The safety prediction problem [5] is defined as follows:
Given horizon $k>0$, a safety predicate $\varphi$, a sequence of $m$ observations and actions from a known a controller $h$: [$\bm{y}_{t},\bm{y}_{t+1},...,\bm{y}_{t+m}$], [$\bm{a}_{t},\bm{a}_{t+1},...,\bm{a}_{t+m}$] from an unknown system $s$, determine the system fits the following logic if the future moments:
@@
Safety prediction is binary, and we use the F1 score to balance precision and recall. Since false-positive safety predictions are dangerous in safety-critical systems, the false positive rate (FPR) is also taken into account.
A major issue with standard world models is that the safety is not straightforward to infer from the observation, so an extra safety evaluator is needed (often a Convolutional Neural Network, CNN) shown on the left of Fig. 2. Our proposed models use meaningful latent states (comparable with the CD metric) where the safety can be checked directly.
In standard world models, the CNN safety evaluator and controller suffer a distribution shift problem because the standard world model cannot perform perfectly and may reconstruct an out-of-distribution observation. The conditional distribution of safety is the same for training and test datasets $p_{train}(\varphi(\bm{x})|\bm{y})=p_{test}(\varphi(\bm{x})|\bm{y})$ but the distribution of observation is shifted: $p_{train}(\bm{y})\neq p_{test}(\bm{y})$, which will influence the accuracy of the CNN evaluator. This issue does not exist in the foundation world models because safety can be interpreted by the estimated states and the image reconstruction is precise pixel movement that will be explained in Sec. III.
@@
Input: A series of segmentation matrices $\Omega(\bm{y})=(\bm{\omega}_{t},\bm{\omega}_{t+1},...,\bm{\omega}_{t+m})$ and corresponding actions $(U=\bm{u}_{t},\bm{u}_{t+1},...,\bm{u}_{t+m})$, and prompt fragment $R$: instruction fragment $r_{1}$, input assembly fragment $r_{2}$, and prediction fragment $r_{3}$; and a state estimation function CentrExt() that extracts the centroid of the object.
@@
Function Assemble($\Omega(\bm{y}),U,R$, CentrExt):
@@
4:     for $j$ from 1 to $\operatorname{len}$($\bm{\omega}_{i}$do
5:         $\pi\leftarrow\pi+r_{2}$triangleright Add assembly prompt
6:         $\bm{z}_{i}^{j}\leftarrow$CentrExt($\bm{\omega}_{i}^{j}$) $\triangleright$ Get predicted state
7:         $\pi\leftarrow\pi+\bm{z}_{i}^{j}$triangleright Add predicted state
8:         $\bm{z}\leftarrow\bm{z}+\bm{z}_{i}^{j}$
9:         $\pi\leftarrow\pi+\bm{u}_{M}$
10:     end $\triangleright$ Add action
@@
12:$\pi\leftarrow\pi+r_{3}$triangleright Add prediction order and specify the required output format
13:return $\pi,\bm{z}$
@@ III Foundation World Models
We propose foundation world models that incorporate (a) a pre-trained foundation segmentation model into our proposed architecture as an encoder and (b) a large language model as a latent predictor. Specifically, we adopt the Segment Anything model (SAM) [7] as the segmentation model to split the observation into several objects: $\Omega(\textit{{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...]$. The extracted centroids $\bm{\omega}^{c}$ are used as latent representations $\bm{z}$, which are causally informative as future object positions are caused (in part) by past object positions. Combining the representations $\bm{z}$ with actions $\bm{u}$ and prompt fragments $R=\{r_{1},r_{2},r_{3}\}$, we adopt the Algorithm 1 to form [^1] We use “+” to denote string concatenation and adding to a state vector. a complete prompt $\pi$. A Large Language Model takes prompt $\pi$ and returns a formatted prediction of ${\bm{P}}(\pi)=\hat{\bm{z}}$. Three types of pre-prompts in $R$ are used to assemble a full prompt: the first $r_{1}$ is to describe the whole task, e.g., “Suppose I have a sequence of actions and states of a system, please predict the next step’s state given the following information.” Next, there’s a loop to fuse the state description $r_{2}$ and the corresponding values “The time, states and the action for this step are:’. We will repeat to add $r_{2}$ for m times to add all the input information. Finally, we add the state prediction instruction $r_{3}$: “Can you predict the state for the next moment? Please only give me your prediction value, and write them into a list” at the end of the prompt to form the $\pi$.
To implement LLM-based prediction of latent states, we choose two models: GPT 3.5 provided by OpenAI and Gemma 7B-it based on the Gemini technology [8]. One challenge is that not all segmented objects prove to be useful; for instance, objects like the white background in the cart pole system are uninformative by themselves. These non-essential objects not only require more tokens in the inference of the Large Language Model, which increases the computational cost and also potentially reduces prediction accuracy. Such redundant information can be eliminated before the step of prompt assembly by removing objects whose centroids do not change at any time in the past observations.
Algorithm 2 Output disassembly into observation and safety label
Input: Predicted state $\bm{z}_{t+m+1}$, observation $\bm{y}_{t+m}$, segmentation matrix $\bm{p}_{t+m}$, and estimated state $\bm{z}_{t+m}$.
Output: Safety label $\varphi$ and predicted observation $\bm{y}_{t+m+1}$
@@
Disassemble ($\bm{z}_{t+m+1},\bm{y}_{t+m},\bm{p}_{t+m},\bm{z}_{t+m}$,SafeEval):
1:$\bm{y}_{t+m+1}\leftarrow\bm{y}_{t+m}$triangleright Copy observation
@@
3:     $\bm{\delta}^{i}\leftarrow\bm{z}_{t+m+1}^{i}-\bm{z}_{t+m}^{i}$triangleright Compute the displacement of each object
4:     for $item=True$ in $\bm{p}_{t+m}$ do
5:         $px\leftarrow item.x$triangleright x position of the pixel
6:         $py\leftarrow item.y$triangleright y position of the pixel
7:         $TEMP\leftarrow\bm{y}_{t+m+1}[px][py]$triangleright store the pixel value temporarily
8:         $\bm{y}_{t+m+1}[px][py]\leftarrow\bm{y}_{t+m+1}[px+{\delta}^{N}_{x}][py+{\delta}^{N}_{y}]$
9:         $\bm{y}_{t+m+1}[px+{\delta}^{N}_{x}][py+{\delta}^{N}_{y}]\leftarrow TEMP$triangleright swap the pixel value
@@
Upright Falling Upright Falling
k=10 k=20 k=30 k=10 k=20 k=30 k=10 k=20 k=30 k=10 k=20 k=30
SAM & GPT 3.5 1.683 3.769 6.398 2.498 6.307 30.50 5.970 12.10 19.37 10.99 23.84 113.8
TABLE I: State prediction performance for the cart pole: the centroid distance for the cart’s position and the mean absolute error (MAE) and the MAE for the pole’s angle. Gray rows indicate the use of additional data.
@@
TABLE II: State prediction performance for the lunar lander: the horizontal and vertical centroid distance errors. Gray rows indicate the use of additional data.
After inputting our prompt $\pi$, we will receive a formatted state prediction output $\hat{\bm{z}}$ based on which we evaluate the safety. To continue the prediction sequence into the next step, we need to generate an observation to feed into the image-based controller $h$. This “output disassembly” step is implemented with Algorithm 2, and describes the process of rebuilding the observation for the controller. This kind of rebuilding won’t cause object duplication and loss as shown in Fig 4, which is a common distribution shift that occurs in standard world models. This problem is brought about by the inability of the decoder to construct an out-of-distribution latent representation predicted by the surrogate model.
Method Input length F1 score FPR

### v2

@@ I Introduction
In the past, researchers primarily focused on training better controllers with world models, with insufficient attention given to how accurate the world of world models is — a crucial aspect for safety-critical robots. Typically, a world model does not directly predict observations, such as images from cameras or other sensor data, due to the constraints on time and computational resources. Instead, as shown on the left of Fig. 2, it extracts useful features into latent representations with an encoder, which is part of a Variational Autoencoder (VAE) [6], and then predicts based on these representations. Usually, the performance of the world model is evaluated with the Mean Square Error (MSE) between the predicted observation and ground truth; this metric considers only the aggregate performance of the reconstruction — and lacks the examination of the fine-grained and meaningful details of the prediction. This coarse checking does not distinguish slight but critical differences (e.g., if a car’s position shifts a few pixels, it may lead to a collision) that may cause devastating outcomes in safety predictions. Also, the latent representations do not have a physical meaning in standard models and cannot be used to evaluate whether a predicted latent state is safe or has other important characteristics. As a result, it becomes necessary to develop an additional classifier to check such critical aspects, akin to a safety check as discussed in our earlier work [5], which requires more data, adds noise, and may suffer from distribution shift in predicted observations (compared to the real sensor data).
@@
Our evaluation on two common simulated benchmarks demonstrates the value of foundation world models. Our foundation model not only demonstrates superior state prediction based on the newly proposed metric but also excels in safety predictions. We experiment with several baseline approaches, evaluation metrics, and different Large Language Models for latent prediction.
@@
    A segmentation-based metric for the accuracy of the surrogate dynamic prediction by quantifying the deviations of each object in the observation.
@@
    An experimental study of safety prediction where foundation world models show better performance despite not using any training data, compared to the existing world model and supervised learning methods.
Sec. II introduces the details of world models and our problem description. Sec. III and Sec. IV describe the foundation world models and the results of our experiments. In Sec. V and Sec. VI, we will review the related work and discuss the conclusion and future work.
@@ II-A Observation Prediction with World Models
There is no single agreed-upon metric to evaluate the performance of a world model. In various applications of world models, the controller-training tasks are usually judged on how well the resulting controller performs compared with traditional methods [1, 3] in terms of reward. Traditionally, evaluation metrics like MSE are commonly adopted as the standard measurement criteria for quantifying the predictive capacity of world models. Accordingly, we compute the MSE of the world model’s predicted observation in a pixel-wise manner:
@@ II-B Object-based Prediction Metric
Our insight, unused in earlier world models, is that the observation can be split into multiple meaningful objects by some segmentation algorithm $\Omega$: $\Omega(\textit{{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...]$. Each $\bm{\omega}$ is a segmentation matrix with $True$ and $False$ pixels with the same size as $\bm{y}$, as shown in Fig. 3 for two example systems. The $True$ elements reflect the position of the segmented object. These objects in the observation are semantically important elements of the world that are assumed to have a causal relation to future positions. The prediction of the objects’ positions will evaluated separately, unlike e.g. the image-wise MSE that takes all pixels into account and is affected by inconsequential entities like static objects/background that occupy most of the observation frame.
To tailor the evaluation to each object $\bm{\omega}$, we will calculate its centroid $\bm{\omega}^{c}=(\bm{\omega}^{c}_{x},\bm{\omega}^{c}_{y})$ (i.e., its center of mass) as:
@@
Each prediction $\hat{\bm{\omega}}$ of a true object $\bm{\omega}$ will be quantified using the centroid distance (CD) defined with some norm $\mathcal{L}$, which can be for instance an L1 or L2-norm:
@@
For evaluating the image-level error between a prediction $\hat{\bm{y}}$ and a ground truth $\bm{y}$, we will implement the segmentation algorithm to split the image into several parts: $\Omega({\bm{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...,{\bm{\omega}}_{n}]$, $\Omega({\hat{\bm{y}}})=[\hat{\bm{\omega}}_{1},\hat{\bm{\omega}}_{2},...,\hat{\bm{\omega}}_{n}]$ get a vector of CDs: $[\text{CD}(\bm{\omega}_{1},\hat{\bm{\omega}}_{1}),\text{CD}(\bm{\omega}_{2},\hat{\bm{\omega}}_{2}),...,\text{CD}(\bm{\omega}_{n},\hat{\bm{\omega}}_{n})]$ and can precisely reflect how well each part is predicted instead of providing all errors of unrelated pixels.
@@ II-C Safety Prediction Problem
Predicate $\varphi$ is a binary function of state $\bm{x}_{t}$ that determines the safety of the system at time $t$: $\varphi(\bm{x}_{t})$. For example, it can indicate if the robot is dangerously close to an obstacle. The safety prediction problem [5] is defined as follows:
Given horizon $k>0$, a safety predicate $\varphi$, a sequence of $m$ observations and actions from a known a controller $h$: [$\bm{y}_{t},\bm{y}_{t+1},...,\bm{y}_{t+m}$], [$\bm{u}_{t},\bm{u}_{t+1},...,\bm{u}_{t+m}$] from an unknown system $s$, determine whether the system satisfies the following logical formula for the future time moments:
@@
Since safety prediction is binary, we use the F1 score to balance the evaluation of precision and recall. False-positive safety predictions are dangerous in safety-critical systems, so the false positive rate (FPR) is also taken into account.
A significant challenge in standard world models is that the safety is not straightforward to infer from the high-dimensional observation, so it necessitates an extra learning-based safety evaluator (often a Convolutional Neural Network, CNN) shown on the left of Fig. 2. This CNN safety evaluator (as well as the controller) can suffer from distribution shift due to distorted predictions $\hat{\bm{y}}$ of standard world models originating in the decoder. Specifically, the conditional distribution of safety is the same for training and test datasets $p_{train}(\varphi(\bm{x})\mid\bm{y})=p_{test}(\varphi(\bm{x})\mid\bm{y})$, but the distribution of observations is shifted: $p_{train}(\bm{y})\neq p_{test}(\bm{y})$, which influences the accuracy of the CNN evaluator.
@@
Input: A series of observation $\bm{y}_{t:t+m}$, a segmentation model $\Omega$ and corresponding actions $(U=\bm{u}_{t},\bm{u}_{t+1},...,\bm{u}_{t+m})$, and prompt fragment $R$: instruction fragment $r_{1}$, input assembly fragment $r_{2}$, and prediction fragment $r_{3}$.
@@
Function Assemble($\bm{y}_{t:t+m},\Omega,U,R$):
@@
4:     $\bm{\omega}_{i}\leftarrow\Omega(\bm{y}_{i})$
5:     for $j$ from 1 to $\operatorname{len}$($\bm{\omega}_{i}$do
6:         $\pi\leftarrow\pi+r_{2}$triangleright Add assembly prompt
7:         $\bm{z}_{i}^{j}\leftarrow$$(\bm{\omega}_{i}^{j})^{c}$triangleright Get the centroid
8:         $\pi\leftarrow\pi+\bm{z}_{i}^{j}$triangleright Add the centroid
9:         $\pi\leftarrow\pi+\bm{u}_{M}$triangleright Add action
10:         $\bm{z}\leftarrow\bm{z}+\bm{z}_{i}^{j}$triangleright Store the centroids for calculating the error later
@@
12:end for
13:$\pi\leftarrow\pi+r_{3}$triangleright Add prediction order and specify the required output format
14:return $\pi,\bm{z}$
@@ III Foundation World Models
We propose foundation world models that incorporate (a) a pre-trained foundation segmentation model into our proposed architecture as an encoder and (b) a large language model as a latent predictor. Specifically, we adopt the Segment Anything model (SAM) [7] as the segmentation model to split the observation into several objects: $\Omega(\textit{{y}})=[\bm{\omega}_{1},\bm{\omega}_{2},...]$. The extracted centroids $\bm{\omega}^{c}$ are used as latent representations $\bm{z}$, which are causally informative as future object positions are caused (in part) by past object positions. Combining the representations $\bm{z}$ with actions $\bm{u}$ and prompt fragments $R=\{r_{1},r_{2},r_{3}\}$, we adopt the Algorithm 1 to form [^1] We use “+” to denote string concatenation and adding to a state vector. a complete prompt $\pi$. A Large Language Model takes prompt $\pi$ and returns a formatted prediction of ${\bm{P}}(\pi)=\hat{\bm{z}}$. Three types of pre-prompts in $R$ are used to assemble a full prompt: the first $r_{1}$ is to describe the whole task, e.g., “Suppose I have a sequence of actions and states of a system, please predict the next step’s state given the following information.” Next, there’s a loop to fuse the state description $r_{2}$ and the corresponding values “The time, states and the action for this step are:’. We will repeat to add $r_{2}$ for m times to add all the input information. Finally, we add the state prediction instruction $r_{3}$: “Can you predict the state for the next moment? Please only give me your prediction values as a list” at the end of the prompt to form the complete $\pi$.
To implement LLM-based prediction of latent states, we choose two models: GPT 3.5 provided by OpenAI and Gemma 7B-it based on the Gemini technology [8]. One challenge is that not all segmented objects prove to be useful; for instance, objects like the white background in the cart pole system are uninformative by themselves. These non-essential objects require more tokens in the inference of the Large Language Model, which increases the computational cost and also potentially reduces prediction accuracy. We eliminate this redundant information before the step of prompt assembly by removing objects whose centroids do not change at any time in the past observations. Future work can explore other causal techniques [9] to focus on the prediction-relevant objects.
Algorithm 2 Output disassembly into observation
Input: Predicted state $\bm{z}_{t+m+1}$, current state $\bm{z}_{t+m}$, current observation $\bm{y}_{t+m}$, and segmentation algorithm $\Omega$.
Output: Predicted observation $\bm{y}_{t+m+1}$
@@
Disassemble ($\bm{z}_{t+m+1},\bm{y}_{t+m},\bm{z}_{t+m},\Omega$):
1:$\bm{y}_{t+m+1}\leftarrow\bm{y}_{t+m}$triangleright Copy most recent observation
@@
3:     $\bm{\omega}_{i}\leftarrow\Omega(\bm{y}_{i})$
4:     $\bm{\delta}^{i}\leftarrow\bm{z}_{t+m+1}^{i}-\bm{z}_{t+m}^{i}$triangleright Compute the displacement of each object
5:     for $pixel=True$ in $\bm{\omega}_{i}$ do
6:         $px\leftarrow pixel.x$triangleright x position of the pixel
7:         $py\leftarrow pixel.y$triangleright y position of the pixel
8:         SWAP($\bm{y}_{t+m+1}[px][py]$, $\bm{y}_{t+m+1}[px+{\delta}^{N}_{x}][py+{\delta}^{N}_{y}]$) $\triangleright$ swap the pixel values
9:     end for
@@
TABLE I: State prediction performance for the lunar lander: the horizontal and vertical centroid distance (CD) errors. Gray rows indicate the use of additional data.
Upright Falling Upright Falling
k=10 k=20 k=30 k=10 k=20 k=30 k=10 k=20 k=30 k=10 k=20 k=30
SAM & GPT 3.5 1.683 3.769 6.398 2.498 6.307 30.50 5.970 12.10 19.37 10.99 23.84 39.76
TABLE II: State prediction performance for the cart pole: the mean centroid distance (CD) and the mean absolute error (MAE) for the pole’s angle. Gray rows indicate the use of additional data.
After inputting our prompt $\pi$, we will receive a formatted state prediction output $\hat{\bm{z}}$ based on which we evaluate the safety. To continue the prediction sequence into the next step, we need to generate an observation to feed into the image-based controller $h$. This “output disassembly” step is implemented with Algorithm 2, and describes the process of rebuilding the observation for the controller. This kind of rebuilding won’t cause object duplication and loss as shown in Fig 4, which is a common distribution shift that occurs in standard world models. This problem is brought about by the inability of the decoder to construct an observation for an out-of-distribution latent representation predicted by the surrogate model. Our proposed models use meaningful latent states, which are comparable based the CD metric, on which safety predicates can be encoded directly and evaluated without learning. This issue does not exist in the foundation world models because safety can be interpreted by the estimated states and the image reconstruction is precise pixel movement.
Method Input length F1 score $\uparrow$ FPR $\downarrow$
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
SAM & MLP .$0$$0$$0$$0$$0$$0$3914 .2845 .2074 .1002 .0535 .0000
SAM & LSTM $m$=1 .$0$$0$$0$$0$$0$$0$0676 .0784 .0579 .0579 .0289 .0000
SAM & MLP .$0$$0$$0$$0$$0$$0$4456 .2692 .2161 .1320 .0400 .0000
SAM & LSTM $m$=2 .$0$$0$$0$$0$$0$$0$0685 .0802 .0536 .0606 .$0$$0$
SAM & Gemma .8850 .9038 .3000 1.000 .8651 .8329 .1441 .$0$$0$$0$0221 .0375
SAM & MLP .$0$$0$$0$$0$$0$$0$5473 .4475 .3860 .3093 .2871 .2889
SAM & LSTM $m$=4 .$0$$0$$0$$0$$0$$0$0825 .0837 .0775 .0597 .$0$$0$
SAM & MLP .$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$
SAM & LSTM $m$=8 .$0$$0$$0$$0$$0$$0$0527 .0667 .0742 .0579 .0460 .0638
SAM & GPT 3.5 .9168 .9710 .6575 .8233 .3260 .6042 .0064 .0068 .$0$$0$$0$$0$
TABLE III: Safety prediction performance for the lunar lander: F1 score and false positive rate. Gray rows indicate the use of additional data.
IV Experiments and Results
IV-A Experimental Setup
We use simulation environments from OpenAI gym [11] as two cases in this paper: (i) a cart pole with a four-dimensional physical state and (ii) a lunar lander with an eight-dimensional physical state. For the cart pole, the safety threshold $\theta_{thre}$ is defined as the angle between the pole and the vertical line $\theta$ being less than $\pi/4$: $|\theta|<\pi/4$. For the lunar lander, safety means keeping the horizontal position $d_{x}$ of the lander within the landing range: $128<d_{x}<$ 12. For the cart pole, to evaluate safety from our predicted states, we calculate the angle instead of showing the pole’s centroid error. For the cart pole, given that the pole is upright most of the time, to get a more detailed evaluation of our models, we split the test experiment of the cart pole into two parts: (i) upright and (ii) falling.
As supervised learning baselines, we train two additional state prediction models: a Multilayer Perceptron (MLP) and a Long Short-Term Memory (LSTM). They are combined with our SAM-based representations. We also compare our approach to the original world model based on VAE representations and an MDN-LSTM state predictor. We test for 10,000 sequences with varied input lengths $m$ and prediction horizons $k$ in each case study. In the results tables, we marked the results of the supervised models in gray because these models take extra 30,000 sequences to train and therefore impose a significant data burden compared to our zero-shot foundation world models. This fact relaxes the expectations of the performance of our proposed models.
Method Input length F1 score $\uparrow$ FPR $\downarrow$